If we sold 10,000 units, had variable costs of $10 per unit, fixed costs of $125,000, and net income of $50,000. What is the selling price per unit?
Answer #1

Let selling price per unit be x

(Contribution margin per unit * units sold) - Fixed cost = Net income

Contribution margin per unit = Selling price per unit - Variable costs per unit

(x - 10)* 10,000 units - 125,000 = 50,000

(x - 10) = (50,000+125,000)/10,000

x - 10 = 17.50

x = 17.50 + 10

Selling price per unit = 27.50